OpenStudy (anonymous):

In how many ways can 1 man and 1 woman, in that order, be selected from 7 men and 8 women?

OpenStudy (anonymous):

7C1 * 8C1

OpenStudy (anonymous):

in that order : \[7\times 8\]

OpenStudy (anonymous):

=56

OpenStudy (anonymous):

is that the formula?

OpenStudy (anonymous):

no that is the 'multiplication principle' which says if i can do one thing n ways and another thing m ways then the number of way i can do them both is n times m

OpenStudy (anonymous):

like those problems that say you have 3 choice of appetizer, 4 of entrées and 2 desserts then you have 3*4*2 different possible meals

OpenStudy (anonymous):

A department store manager wants to display 7 brands of a product along one shelf of an aisle. In how many ways can he arrange the brands? so you would do 7*1?

OpenStudy (anonymous):

think that you have seven slots in which to put 7 items. you have 7 choices for the first one. so far 7 ways

OpenStudy (anonymous):

now one brand is used and so you have 6 choices for the next slot. so far 7*6

OpenStudy (anonymous):

now you have 5 choices for the fifth slot, 4 choices for the 4th etc. and you have to multiply all these together by the multiplication principle

OpenStudy (anonymous):

answer is \[7\times 6\times 5\times 4\times 3\times 2 \times 1\]

OpenStudy (anonymous):

this number is also written as 7!

OpenStudy (anonymous):

ur awesome lol

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

i have a worksheet on scp im reading the textbook but it dosent break it down for me to good so im just trying to understand it

OpenStudy (anonymous):

scp?

OpenStudy (anonymous):

sequential counting principle

OpenStudy (anonymous):

got it any more ?

OpenStudy (anonymous):

o yea lol

OpenStudy (anonymous):

got time for one more if you want to ask

OpenStudy (anonymous):

At a high school cafeteria, diners can choose one vegetable from a choice of 4 vegetables, one meat from a choice of 3 meats, one serving of bread from among 4 breads, and a dessert from among 2 desserts. How many meal configurations are possible?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

we just did this remember? look above!

OpenStudy (anonymous):

i would do 4*3*4*2?

OpenStudy (anonymous):

xactly

OpenStudy (anonymous):

yes but what confused me is when it says choose one vegetable from a choice of 4 vegetables,

OpenStudy (anonymous):

so i thougt i had to do the 4*3*2*1

OpenStudy (anonymous):

yes one choice of the 4, so you have 4 choices. silly english i know, but that is what it means.

OpenStudy (anonymous):

now i will go have lunch myself. you look to be in good shape for this.

OpenStudy (anonymous):

have a good lunch.. wish i can buy for helping me